# From South to North: Leveraging Ground-Based LATs for Full-Sky CMB Delensing and Constraints on $r$

### v1

@@ 3.2 Needlet Internal Linear Combination
$\hat{X}_{\text{NILC}}(\hat{\mathbf{n}})=\sum_{\ell m}\sum_{jk}\beta_{jk}^{\mathrm{NILC}}\sqrt{\frac{4\pi}{\lambda_{j}}}h_{\ell}^{j}Y_{\ell m}(\xi_{j})Y_{\ell m}(\hat{\mathbf{n}}).$ (3.9)
@@ 3.3 Map combination
We emphasize that this work employs the diagonal filtering approach, despite its known sub-optimal performance for cut-sky analyses when compared to the more sophisticated methodology described in [51]. This decision is based on two considerations. First, we have tested the alternative method and found it to be computationally demanding. Second, we do not expect a significant improvement in delensing efficiency from this adjustment. As described by [52], the key factor determining delensing efficiency is the accuracy of the lensing estimate, rather than contamination in the CMB maps undergoing delensing.
In this paper, we first combine the NILC-cleaned maps from the two LATs (hereafter denoted as LATN+LATS), both of which cover a multipole range of (0,6143)(0,6143)( 0 , 6143 ), over the sky regions shown in Fig. 2. The outputs are the combined $\Theta$, $E$, and $B$ maps. These maps serve two purposes. First, they are used for internal lensing reconstruction. Second, they are further combined with the NILC-cleaned maps from the SAT (hereafter denoted as LATN+LATS+SAT), where the combined LAT covers the multipole range (0,6143)(0,6143)( 0 , 6143 ), and the SAT covers (0,300)(0,300)( 0 , 300 ). The final combined maps are then utilized for lensing B-mode template construction and parameter constraints. For comparative analysis, we additionally generate a LATS+SAT case by applying identical combination procedures to the NILC-cleaned maps from the LAT South (LATS) and SAT instruments.
@@ 3.4.1 Internal Lensing Reconstruction
For polarization reconstruction, we use a multipole range of (200,5600)(200,5600)( 200 , 5600 ), excluding multipoles below 200 to mitigate biases introduced in the delensing procedure. These biases arise from the overlap between the B-modes used for reconstructing $\phi$ and the B-modes intended for delensing (i.e., the observed B-modes), as noted by [53, 54]. For temperature reconstruction, we restrict the analysis to multipoles in the range (200,5200)(200,5200)( 200 , 5200 ) to relieve contamination from large-scale atmospheric noise and residual small-scale extragalactic foreground contributions, as discussed in [55, 56].
@@
${{}_{1}}\hat{d}(\hat{\mathbf{n}})=-\sum_{s=0,\pm 2}{{}_{-s}}\bar{X}(\hat{\mathbf{n}})\big{[}\sharp_{s}X^{\text{WF}}\big{]}(\hat{\mathbf{n}}).$ (3.11)
Here, $X=\{\Theta,E,B\}$ from observation, $\sharp$ denotes the spin-raising operator, which is associated with the gradient of a field [27], and the subscript $s$ represents the spin of the fields. The estimator takes as input the gradient of the Wiener-filtered maps, ${}_{s}X^{\text{WF}}(\hat{\mathbf{n}})$, and the real-space inverse-variance filtered maps, ${}_{s}\bar{X}(\hat{\mathbf{n}})$.
For filtering, we apply the optimal filtering method described in [57], implemented using the multi-grid preconditioned conjugate gradient algorithm. The iteration tolerance level $\epsilon$ is set to $10^{-5}$ for temperature filtering and $10^{-4}$ for polarization filtering, which is considered sufficiently precise for accurate lensing reconstruction. The root-mean-square (RMS) map required by this method is chosen to be an isotropic noise RMS map (we do not incorporate the hit-count variations in this work), with effective noise levels given in Table 3. These noise levels are determined from the power spectra of the null signal simulation, $X^{\text{NILC,null},i}$, for each experiment, thereby accounting for contributions from both instrumental noise and foreground residuals:
$\hat{\mathbf{n}}_{\text{lev}}=\sqrt{\langle\hat{C}_{\ell}^{\text{noise}}\rangle_{\ell}}\frac{60\cdot 180}{\pi},$ (3.12)
where $\langle\rangle_{\ell}$ denotes the average over the multipole range $2200<\ell<2600$ for temperature and $700<\ell<1000$ for polarization. This range is not strictly defined and is somewhat empirical, but it is chosen to suppress noisy modes at the smallest scales, where lensing reconstruction primarily relies on CMB information, as suggested in [29]. Additionally, we apply an apodized mask to reduce the creation of spurious gradients at the mask boundary [58] with an apodization scale of $6^{\prime}$, which has been shown to significantly accelerate the convergence of the iterative filtering process, particularly for temperature filtering.
Table 3: The effective white noise level used for lensing reconstruction filtering includes contributions from both instrumental noise and foreground residuals. Notably, in the overlapping regions, the polarization noise decreases by nearly $1/\sqrt{2}$ as expected, whereas the temperature noise shows only a slight reduction. We attribute this mainly to the large-scale foreground residuals, which are highly correlated since only one single foreground realization was used in the mock data.
nlev ($\mu$K-arcmin) LATN exclusive regions LATS exclusive regions Overlapping regions
Temperature 16.7 16.8 14.6
@@
[…]
@@ 3.4.2 External Lensing Reconstruction
For the CIB, we use only the simulated maps at 353 GHz, where the GNILC CIB map has been shown to exhibit the highest correlation with CMB lensing [59]. Moreover, given the high signal-to-noise ratio (S/N) of the CIB auto-power spectrum over the relevant angular scales, the contribution of instrumental noise is negligible. Consequently, we omit instrumental noise in our theoretical modeling of the CIB.
For galaxy number density, we adopt the configuration of the Euclid survey, which achieves an effective galaxy number density of approximately 30 arcmin-2 [33]. To fully exploit the redshift information provided by Euclid, we divide the galaxy sample into five tomographic redshift bins, each serving as a tracer of the underlying matter distribution. It is worth noting that we assume the Euclid survey fully overlaps with the sky coverage of the ground-LATs’ configuration. In practice, some deviation is expected due to the differing scan strategies and scientific objectives of the two surveys, we leave a more detailed study to future work.
@@ 3.5.1 Constructing LT on map-level
In this paper, we construct the lensing B-mode template using two delensing methods, as discussed in Section 2. To assess the improvement gained by incorporating LATN data and LSS tracers, we construct the lensing $B$-mode template (LT) separately for four cases: LATS+SAT, LATS+SAT+External, LATN+LATS+SAT, and LATN+LATS+SAT+External. In each case, the LT is constructed only within the overlapping region between the ground-based LATs and SAT. This restriction is not only due to the lack of lensing potential estimates in the non-overlapping regions, but also because of the absence of smaller-scale ($\ell>300$) lensed $E$-mode observations (see FIG.2 of [32]), both of which are essential for accurate LT construction.
@@
$\begin{gathered}Q_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}Q_{\ell m},\\ U_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}U_{\ell m},\\ \end{gathered}$ (3.16)
@@
Regarding the Gradient-order lensing template method, we proceed by converting the spin-2 fields $Q+iU$ to spin-1 and spin-3 fields, and transform the spin-0 lensing potential field to spin-1-1- 1 and spin-1 fields using CMBlensplus [68]. Following this conversion, we compute the gradient lensing template according to Section 2. The resulting QU represents the lensing effect on E-mode, from which the lensing B-mode can be separated using harmonic transformation and EB leakage is corrected with an ”E-mode Recycling” method described in [49].
@@ 3.5.2 Bandpower calculation and bias estimation
We compute the auto- and cross-bandpowers of lensing B-mode template (LT) and foreground-cleaned observed B-mode ($C_{b}^{\text{LT}}$ and $C_{b}^{\text{LT}\times\text{obs}}$) using the NaMaster[50] code. The multipole range for likelihood analysis (Section 3.6) spans from $\ell_{\text{min}}=20$ to $\ell_{\text{max}}=200$, with a binning width of $\Delta_{\ell}=20$.
As for the corresponding bias terms analyzed in Section 2.2, we input the required CMB maps and the potential maps to the same pipeline as described above to simulate all the bias terms, their auto- and cross-bandpowers are then also calculated using the NaMaster code. As mentioned in Section 2.2, we can also replace the reconstructed $\phi$ with the true $\phi$ in the baseline, which will cause some of the cross terms in the results to vanish. This allows us to isolate these terms by taking the difference between the baseline simulation results and these specific simulation results. We found that these two debiasing methods have similar effects, allowing us to choose either one flexibly. The average over 500 simulations of these terms serves as an estimate of the bias terms in the likelihood.
@@ 3.6 Likelihood analysis and $r$ constraint
The priors on the parameters are imposed as uniform distributions, as listed in Table 4. We assume a Gaussian likelihood to obtain the posterior distribution of parameters, which is a reasonable approximation given the large sky patch used in our analysis. The Gaussian approximation for the log-likelihood function adopted in this work is given by:
$\begin{gathered}-2\ \text{ln}\ \mathcal{L}(C_{b}|\hat{C_{b}})=(\mathbf{X}_{b}-\mathbf{\hat{X}}_{b})^{T}\mathbf{M_{b}^{-1}}(\mathbf{X}_{b}-\mathbf{\hat{X}}_{b})+\text{ln}\ |\mathbf{M_{b}}|,\end{gathered}$ (3.22)
where the vector $\mathbf{\hat{X}}_{b}$ contains the band-power estimates derived from the observed cut-sky maps, computed using the NaMaster code.
@@ 4.1 Results of lensing reconstruction
Figure 6: Reconstructed lensing map for LATN + LATS case. The left panel is the input lensing map; the middle panel is the reconstructed map from the MV quadratic estimator, the right panel is the difference between them. In order to highlight the lens structure, we plot the Wiener -filtered lensing deflection angle amplitude $\hat{\alpha}^{\text{WF}}_{LM}=\sqrt{L(L+1)}\hat{\phi}_{LM}^{\text{MV}}C^{\phi\phi,\text{fid}}_{L}/[C^{\phi\phi,\text{fid}}_{L}+N^{(0),\text{ana}}_{L}]$.
@@
Figure 7: The lensing power spectra from the internal reconstruction (MV) of a single simulation for both the LATS and LATN+LATS cases are shown. The light-blue and gold shaded regions represent the $1\sigma$ standard deviation estimated from 300 simulations for the LATS and LATN+LATS cases, respectively. The terms ”$N^{0}_{L}$” denote the semi-analytical leading-order reconstruction noise, which is used for Wiener filtering.
Figure 8: The correlation coefficient $\rho$ of each tracer, as well as the combined tracer (black), with the true convergence is shown. It is worth noting that improvements in $\rho$ at different angular scales contribute unequally to the effectiveness of delensing. In particular, the inclusion of external tracers significantly enhances the correlation on smaller scales compared to internal lensing reconstruction alone, which can lead to a non-negligible improvement in delensing performance.
@@ 4.2 Results of LT construction
[…]
@@
Figure 10: The power spectra of the NILC-cleaned B-modes and the lensing B-mode template (LT) are shown for two cases, with all spectra properly debiased. The LT was generated using the gradient-order method, and the power spectra have been debiased and rescaled using the transfer functions. The black lines represent the theoretical predictions based on our models with best-fit parameters. In the top panel, the blue and green shaded areas indicate the $1\sigma$ standard deviation computed from 500 observation simulations for the LATN+LATS+SAT(same as LATN+LATS+SAT+External) and LATS+SAT(same as LATS + SAT + External) cases, respectively. In the lower two panels, the red, yellow and blue shaded areas represent the $1\sigma$ standard deviation of the LT for the LATN+LATS+SAT+External, LATN+LATS+SAT and LATS+SAT cases, respectively. As observed, all mean spectra in the lower two panels, after debiasing, align well with the theoretical predictions, which are precisely overlaid by the other curves. Moreover, a clear reduction in uncertainty is evident when comparing the LT derived from the LATN+LATS+SAT case to that from the LATS+SAT case.

### v2

@@ 3.2 Needlet Internal Linear Combination
[…]
@@ 3.3 Map combination
We emphasize that our analysis employs a diagonal inverse-variance weighting in harmonic space for map combination, together with the Wiener filtering applied in harmonic space in the LT construction (Section 3.5). Taken together, these two steps yield a sub-optimal performance for cut-sky analyses compared to a more sophisticated methodology described in [42], which may be regarded as a more complete Wiener-filtered treatment of the observations. This decision is based on two considerations. First, we have tested the alternative method and found it to be computationally demanding. Second, we do not expect a significant improvement in delensing efficiency from this adjustment. As described by [75], the key factor determining delensing efficiency is the accuracy of the lensing estimate, rather than contamination in the CMB maps undergoing delensing.
In this paper, we first combine the NILC-cleaned maps from the two LATs (hereafter denoted as LATN+LATS), both of which cover a multipole range of $(0,6143)$, over the sky regions shown in Fig. 2. The outputs are the combined $\Theta$, $E$, and $B$ maps. These maps serve two purposes. First, they are used for internal lensing reconstruction. Second, they are further combined with the NILC-cleaned maps from the SAT (hereafter denoted as LATN+LATS+SAT), where the combined LAT covers the multipole range $(0,6143)$, and the SAT covers $(0,300)$. The final combined maps are then utilized for lensing B-mode template construction and parameter constraints. For comparative analysis, we additionally generate a LATS+SAT case by applying identical combination procedures to the NILC-cleaned maps from the LAT South (LATS) and SAT instruments.
@@ 3.4.1 Internal Lensing Reconstruction
For polarization reconstruction, we use a multipole range of $(200,5200)$, excluding multipoles below 200 to mitigate biases introduced in the delensing procedure. These biases arise from the overlap between the B-modes used for reconstructing $\phi$ and the B-modes intended for delensing (i.e., the observed B-modes), as noted by [72, 112]. For temperature reconstruction, we restrict the analysis to multipoles in the range $(200,5200)$ to relieve contamination from large-scale atmospheric noise and residual small-scale extragalactic foreground contributions, as discussed in [120, 17].
@@
[…]
For filtering, we apply the optimal filtering method described in [79], implemented using the multi-grid preconditioned conjugate gradient algorithm. The iteration tolerance level $\epsilon$ is set to $10^{-5}$ for both temperature filtering and polarization filtering, which is considered sufficiently precise for accurate lensing reconstruction. The root-mean-square (RMS) map required by this method is chosen to be an isotropic noise RMS map (we do not incorporate the hit-count variations in this work), with effective noise levels given in Table 3. These noise levels are approximately determined from the power spectra of the null signal simulation, $X^{\text{NILC,null},i}$, for each experiment, thereby accounting for contributions from both instrumental noise and foreground residuals:
$\hat{\mathbf{n}}_{\text{lev}}=\sqrt{\sum_{\ell}\frac{2\ell+1}{4\pi}\hat{C}_{\ell}^{\text{noise}}\cdot\Omega_{\text{pix}}}\frac{60\cdot 180}{\pi}[\mu\text{K-arcmin}],$ (3.12)
where the summation range is identical to that used in the lensing reconstruction, $\hat{C}_{\ell}^{\mathrm{noise}}$ denotes the power spectrum of the null-signal simulation $X^{\mathrm{NILC,null},i}$, and $\Omega{\mathrm{pix}}=\frac{4\pi}{12\ast\mathrm{NSIDE}^{2}}$ is the pixel area of a HEALPix map. Additionally, we apply an apodized mask to reduce the creation of spurious gradients at the mask boundary [82] with an apodization scale of $6^{\prime}$, which has been shown to significantly accelerate the convergence of the iterative filtering process, particularly for temperature filtering.
Table 3: The effective white noise level used for lensing reconstruction filtering includes contributions from both instrumental noise and foreground residuals. As expected, in the overlapping regions, both the temperature and the polarization noise decreases by nearly $1/\sqrt{2}$ due to the multiple realizations of both instrumental noise and foreground.
[…]
Temperature 20.0 18.6 13.6
@@
where the gradient mode $\hat{g}_{LM}$ is obtained from the quadratic terms of the inverse-filtered and Wiener-filtered fields. The response function $\mathcal{R}_{L}^{\phi}$ serves as a normalization factor and must be computed accordingly. We apply an approximate normalization at the map level, calculated analytically following [86]. Although this semi-analytic normalization is only accurate within $1\%-2\%$, it is sufficient for the purpose of map-level delensing.
@@ 3.4.2 External Lensing Reconstruction
For the CIB, we use only the simulated maps at 353 GHz, where the GNILC CIB map has been shown to exhibit the highest correlation with CMB lensing [122]. Moreover, given the high signal-to-noise ratio (S/N) of the CIB auto-power spectrum over the relevant angular scales, the contribution of instrumental noise is negligible. Consequently, we omit instrumental noise in our theoretical modeling of the CIB. We assume the CIB fully overlaps with the sky coverage of the SAT’s configuration in this work, and the large-scale CIB ($L<100$) are excluded to avoid the contamination from Galactic foreground residuals [80].
For galaxy number density, we adopt the configuration of the Euclid survey, which achieves an effective galaxy number density of approximately 30 arcmin-2 [65]. To fully exploit the redshift information provided by Euclid, we divide the galaxy sample into five tomographic redshift bins, each serving as a tracer of the underlying matter distribution. It is worth noting that we assume the Euclid survey fully overlaps with the sky coverage of the SAT’s configuration. In practice, some deviation is expected due to the differing scan strategies and scientific objectives of the two surveys. However, thanks to the forthcoming fourth- and fifth-generation galaxy surveys—including DESI [1], LSST [58], the Nancy Grace Roman Space Telescope [107], Euclid [78], CSST [46], MUST [123], the MegaMapper telescope [99], the 11 m Maunakea Spectroscopic Explorer (MSE) [14], the 12 m Wide-field Spectroscopic Telescope (WST) [103], and the $\sim$12 m Extremely Large Spectroscopic Survey Telescope (ESST) [110]—we expect that combining these surveys will provide a deep and wide galaxy sample suitable for realistic data processing. A more detailed investigation is left to future work.
@@ 3.5.1 Constructing LT on map-level
In this paper, we construct the lensing B-mode template using two delensing methods, as discussed in Section 2. To assess the improvement gained by incorporating LATN data and LSS tracers, we construct the lensing $B$-mode template (LT) separately for four cases: LATS+SAT, LATS+SAT+External, LATN+LATS+SAT, and LATN+LATS+SAT+External. In the internal delensing cases (LATS+SAT and LATN+LATS+SAT), the LT is constructed only within the sky region overlapping between the ground-based LATs and the SAT, since the lensing field is not available outside this area. In the joint delensing cases (LATS+SAT+External and LATN+LATS+SAT+External), the LT can in principle be constructed over the full sky; however, for a consistent and fair baseline comparison, we restrict the analysis to the same overlapping sky region used in the internal delensing scenarios. Results that additionally include the LT outside the LAT–SAT overlapping area are presented in the Appendix C.
@@
[…]
@@
Regarding the Gradient-order lensing template method, we proceed by converting the spin-2 fields $Q+iU$ to spin-1 and spin-3 fields, and transform the spin-0 lensing potential field to spin-1-1 and spin-1 fields using CMBlensplus [85]. Following this conversion, we compute the gradient lensing template according to Section 2. The resulting QU represents the lensing effect on E-mode, from which the lensing B-mode can be separated using harmonic transformation and EB leakage is corrected with an ”E-mode Recycling” method described in [70].
@@ 3.5.2 Bandpower calculation and bias estimation
We compute the auto- and cross-bandpowers of lensing B-mode template (LT) and foreground-cleaned observed B-mode ($C_{b}^{\text{LT}}$ and $C_{b}^{\text{LT}\times\text{obs}}$) using the NaMaster[12] code. The multipole range for likelihood analysis (Section 3.6) spans from $\ell_{\text{min}}=2$ to $\ell_{\text{max}}=200$, with a binning width of $\Delta\ell=40$.
As discussed in Sec. 3.5, the lensing template (LT) is defined only over the LAT–SAT overlap region, whereas the observed B-mode map covers the full survey footprint after applying the Galactic mask. For the auto-spectra, we apply apodized masks independently to each field (with a 0.5° apodization for the LT and 0.2° for the observed B modes, generated using the NaMaster Smooth function). In contrast, when computing the cross-spectrum between the LT and the observed B modes, the mask must correspond to the intersection of their sky coverages. Only this common region contains valid LT information, and therefore only this region provides a physically meaningful contribution to the cross-spectrum. The resulting bandpowers are then included in the data vector (Eq. 3.18) used for the likelihood analysis.
As for the corresponding bias terms analyzed in Section 2.2, we input the required CMB maps and the potential maps to the same pipeline as described above to simulate all the bias terms, their auto- and cross-bandpowers are then also calculated using the NaMaster code. An alternative method is to replace the reconstructed $\phi$ with the true $\phi$ in the baseline, which will cause some of the cross terms in the results to vanish. This allows us to isolate these terms by taking the difference between the baseline simulation results and these specific simulation results. We found that these two debiasing methods have similar effects, allowing us to choose either one flexibly. The average over 500 simulations of these terms serves as an estimate of the bias terms in the likelihood.
@@ 3.6 Likelihood analysis and $r$ constraint
Considering that our analysis includes the reionization $B$-modes at low multipoles (where only a small number of modes are available), we adopt the Hamimeche–Lewis (HL) likelihood [50] in the baseline likelihood analysis. The HL likelihood effectively captures the non-Gaussian distribution of bandpowers in this low-$\ell$ regime and thus provides a more accurate treatment than a simple Gaussian approximation:
$-2\ln\mathcal{L}(C_{b}|\hat{C}_{b})=\frac{2b+1}{2}\mathrm{Tr}\left[\left(\mathbf{C}_{f_{b}}^{-1/2}\mathbf{C}_{g_{b}}\mathbf{C}_{f_{b}}^{-1/2}\right)^{2}\right]=\mathbf{X}_{g_{b}}^{T}\mathbf{M_{b}}^{-1}\mathbf{X}_{g_{b}},$ (3.22)
where $\mathbf{X}_{g_{b}}=\text{vecp}(C_{g_{b}})$ is the vector of distinct elements of the transformed power spectra $C_{gb}$. For a comprehensive review, we refer the reader to reference [50]. A discussion about the alternative choice of the likelihood function is given in Appendix B.2.
@@
Since the covariance is estimated from a finite number of simulations, appropriate corrections are required to ensure an accurate likelihood analysis, as implemented in many previous works [23, 36, 7, 2, 5, 18, 33, 41]. We first apply the Hartlap–Anderson factor [52] to the inverse covariance matrix, $\mathbf{M}_{b}^{-1}$, to correct for the bias introduced by Monte-Carlo noise in the covariance estimate. In addition, constraining noisy elements of the covariance matrix (i.e., by conditioning the covariance estimate) is essential to reduce errors arising from the finite number of realizations. Our choice of covariance-conditioning scheme is described in Appendix B.1.
@@ 4.1 Results of lensing reconstruction
Figure 6: Reconstructed lensing map for the LATN+LATS case. The upper-left panel shows the input lensing map, and the lower-left panel shows the reconstructed map using the MV quadratic estimator (the overlap region between LATN and LATS is outlined in the lower-right panel). The upper-right panel shows the difference between the two maps. To better highlight the lensing structure, we plot the Wiener-filtered deflection-angle amplitude $\hat{\alpha}^{\mathrm{WF}}_{LM}=\sqrt{L(L+1)}\,\hat{\phi}^{\mathrm{MV}}_{LM}\,\frac{C^{\phi\phi,\mathrm{fid}}_{L}}{C^{\phi\phi,\mathrm{fid}}_{L}+N^{(0),\mathrm{ana}}_{L}}$. A clear improvement in signal-to-noise ratio is observed in the overlap region between LATN and LATS.
@@
Figure 7: The lensing power spectra from the internal reconstruction (MV) of a single simulation for both the LATS and LATN+LATS cases are shown. The light-blue and gold shaded regions represent the $1\sigma$ standard deviation estimated from 400 simulations for the LATS and LATN+LATS cases, respectively. The terms $N^{(0)}_{L}$” denote the semi-analytical leading-order reconstruction noise, which is used for Wiener filtering.
Figure 8: The correlation coefficient $\rho$ of each tracer, as well as the combined tracer (black), with the true convergence is shown. Here ”CMB” represents the internal reconstruction for LATN+LATS case. It is worth noting that improvements in $\rho$ at different angular scales contribute unequally to the effectiveness of delensing. In particular, the inclusion of external tracers significantly enhances the correlation on smaller scales compared to internal lensing reconstruction alone, which can lead to a non-negligible improvement in delensing performance.
@@ 4.2 Results of LT construction
Figure 9: The delensed $B$-mode maps obtained using the gradient-order method for both the LATN+LATS+SAT and LATS+SAT cases. The two middle zoomed-in subplots depict a $10^{\circ}\times 10^{\circ}$ region centered at $(50^{\circ},70^{\circ})$ in Galactic coordinates. Shaded regions indicate the lensing $B$-mode reduction. Compared to LATS+SAT ($43\%$ sky coverage), the inclusion of LATN data expands coverage to $61\%$, improving the lensing template reconstruction.
@@
Figure 10: The power spectra of the NILC-cleaned B-modes and the lensing B-mode template (LT) are shown for two cases, with all spectra properly debiased. The LT was generated using the gradient-order method, and the power spectra have been debiased and rescaled using the transfer functions. The black lines represent the theoretical predictions based on our models with best-fit parameters. In the top panel, the blue and green shaded areas indicate the $3\sigma$ standard deviation computed from 500 observation simulations for the LATN+LATS+SAT(same as LATN+LATS+SAT+External) and LATS+SAT(same as LATS + SAT + External) cases, respectively. In the lower two panels, the red, yellow and blue shaded areas represent the $2\sigma$ standard deviation of the LT for the LATN+LATS+SAT+External, LATN+LATS+SAT and LATS+SAT cases, respectively. As observed, all mean spectra in the lower two panels, after debiasing, align well with the theoretical predictions, which are precisely overlaid by the other curves. Moreover, a clear reduction in uncertainty is evident when comparing the LT derived from the LATN+LATS+SAT case to that from the LATS+SAT case.